write the equation of the line through the given points. through (2,0) and (-2,1)
Answer by Alan3354(69443)

write the equation of the line through the given points. through (2,0) and (-2,1)
------------------
Step 1, find the slope, m
--
m = diffy/diffx = (1-0)/(-2-2) = -1/4
================
Step 2, use y - y1 = m*(x - x1) where (x1,y1) is either point.
=====================
PS, not "the equation"
An equation.